OpenStudy (anonymous):

Please teach me how to find the focal length of a convex lens by using displacement method :(

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

yea...wat u do is, try to focus a distance light source say sun(highly prefered) and try to get them converge onto a single point. or atleast try to focus such that the disk formed on the screen is the smallest, then measure the distance between the lens and the screen, u hav ur focal length.

OpenStudy (anonymous):

We have to use uprights (needles) to find the focal length. How do we do it by them?

OpenStudy (anonymous):

We find the approximate f by the sun. And then adjust the position of object and image needles to find this one point. :/

OpenStudy (anonymous):

this displacement method is buried somewher in my physics register

OpenStudy (anonymous):

keep a convex lens in front of a screen.. then keep a fixed upright at distance D now move the lens along the line 1/ (D-x) - 1/(-x) = 1/f where x is the distance of the lens from the fixed upright object... so D-x in +ve direction is the image distance..because were taking x to be any position for which an image is formed on the screen So u get a 2 deg eqn in x x^2 - Dx +Df = 0 Ull get 2 values of x for which an image is formed..u can measure the distance bw these 2 points using a metre scale on the optical bench..lets say the measured distance is t t = x2 - x1 (difference of 2 x-values) t = root(D^2 - 4Df) U know D and t, so find f :)

OpenStudy (anonymous):

Oh right. Got it. Thanks.

OpenStudy (anonymous):

u made sure i learnt it as well...somewhere deep in my optics copy